# Simple-RF: Regularizing Sparse Input Radiance Fields with Simpler Solutions

### v4

@@ 7.2. Limitations and Future Work
Our approach of obtaining reliable depth supervision by learning simpler solutions is limited to the cases where the base model achieves a reasonable performance but suffers from various distortions due to overfitting. Our regularizations may not help significantly if the performance of the base model is very poor, as in the case of learning a highly complex scene with very few input views as shown in Fig. 27. In such cases, we can employ our regularizations on top of other sparse input radiance fields as we do for the NeRF-LLFF dataset.
@@
Finally, our approach of employing augmentations to obtain better supervision can be explored in sparse input novel view synthesis of highly specular objects (Verbin et al., 2022), refractive objects (Deng et al., 2024), low-light scenes (Mildenhall et al., 2022), blurred images (Ma et al., 2022) and high dynamic range images (Lu et al., 2024). Further, it would be interesting to explore the use of augmentations in related inverse problems such as surface reconstruction (Wang et al., 2021a; Long et al., 2022), dynamic view synthesis (Pumarola et al., 2021; Fridovich-Keil et al., 2023; Somraj et al., 2024, 2022; Choudhary et al., 2025) and style transfer (Wang et al., 2024a) when only sparse input viewpoints are available.

### v5

@@ 7.3. Limitations and Future Work
Our approach of obtaining reliable depth supervision by learning simpler solutions is limited to the cases where the base model achieves a reasonable performance but suffers from various distortions due to overfitting. Our regularizations may not help significantly if the performance of the base model is very poor, as in the case of learning a highly complex scene with very few input views as shown in Fig. 27. In such cases, we can employ our regularizations on top of other sparse input radiance fields as we do for the NeRF-LLFF dataset. Sec. 7.2 provides a couple of approaches to determine if our model is likely to help improve the performance for a given scene.
Even though our method significantly reduces the floaters, some translucent floaters still persist as seen in the supplementary videos. Our approach provides supervision from the augmented models only at the pixels where reliable depth estimates are available. We determine the reliable depths by reprojecting to the closest training view and checking for consistency as explained in Sec. 4.1.4. Hence, in regions which are visible in only one training view, we do not have reliable depth supervision, which sometimes leads to transparent floaters. Further, we observe significant transparent floaters in the TensoRF model. Although we employ a mass concentration loss (Eq. 15) to increase opacity, we observe that the transparent objects are not completely removed. We also note that employing this loss in other models did not help in removing transparent floaters. Finally, our depth supervision loss is imposed only on the mean depth predicted by the model (Eq. 9) and not on the variance of the distribution. We experimented with imposing an additional loss to reduce transparent floaters by minimizing the standard deviation of the predicted depth or a KL divergence loss between the predicted depth distribution and a delta function at the estimated depth prior (Deng et al., 2022), we found that it led to several stronger artifacts. This perhaps requires a more sophisticated approach and is left for future work.
@@
Finally, our approach of employing augmentations to obtain better supervision can be explored in sparse input novel view synthesis of highly specular objects (Verbin et al., 2022), refractive objects (Deng et al., 2024), low-light scenes (Mildenhall et al., 2022), blurred images (Ma et al., 2022) and high dynamic range images (Lu et al., 2024). Further, it would be interesting to explore the use of augmentations in related inverse problems such as surface reconstruction (Wang et al., 2021a; Long et al., 2022), dynamic view synthesis (Pumarola et al., 2021; Fridovich-Keil et al., 2023; Somraj et al., 2024, 2022; Choudhary et al., 2025) and style transfer (Wang et al., 2024b) when only sparse input viewpoints are available.